Calculate the optical invariant for a photographic objective with F = 50 mm. N = 2, field of view +26.6°.


1
Expert's answer
2021-04-05T11:10:06-0400

Given,

"F=50mm=50\times 10^{-3}m"

"N=2"

View angle "(\theta) =\pm 26.6^\circ"

Optical invariant "(I)=FN\sin\theta"

Now, substituting the values,

"(I)=50\times 10^{-3}\times 2\times \sin(26.6^\circ)"

"\Rightarrow (I)=0.1\times \sin(26.6^\circ)"

"=0.1\times 0.447"

"=0.0447"